determine the volume of .0100M NaOH needed to titrate 20mL of .400M HCl


1
Expert's answer
2021-02-18T07:33:14-0500

NaOH+HCl=NaCl+H2O

n (NaOH)=n(HCl)=CM x V

0.0100 x V (NaOH) = 0.400 x 20

V (NaOH) = (0.400 x 20)/0.0100 = 800 ml